OpenStudy (anonymous):

explain how you would fine the exact value of 3/sqrt8

OpenStudy (ybarrap):

set = x and complete the square

OpenStudy (anonymous):

i tried it and got 3*sqrt2/4

OpenStudy (ybarrap):

3/sqrt(8) = x x^2 = 9/8 x^2 -9/8=0 (x + sqrt(9/8))(x-sqrt(9/8)) =0

OpenStudy (queelius):

Well, what you have given is already an exact value. Are you asking for some sort of simplification? See if you can follow this? \[\frac{ 3 }{ \sqrt{8} } = \frac{ 3 }{ \sqrt{2^3} } = \frac{ 3 }{ 2\sqrt{2} } = \frac{ 3\sqrt{2} }{ 4 }\]

OpenStudy (anonymous):

yea thats what i got @queelius

OpenStudy (anonymous):

Or, you could replace 3 with the square root of 9: \[\frac{ \sqrt{9} }{ \sqrt{8} } = \sqrt{9/8}\]But as mentioned above, your original expression was already exact.

OpenStudy (anonymous):

idk wat the correct answer is

OpenStudy (queelius):

Could you repeat the exact question you're trying to answer?

OpenStudy (anonymous):

\[3/\sqrt8\] explain how to find the exact value

OpenStudy (queelius):

That is a confusing question. It would make more sense if they asked for an approximate value. :) I think what I posted, which is what you had previously arrived at, is the expected answer.

OpenStudy (ybarrap):

I think you are asking how do you find the decimal representation of 3/sqrt(8)

OpenStudy (anonymous):

that is what i thought because its a short response

OpenStudy (anonymous):

no @ybarrap it actuallu has a square root form

OpenStudy (anonymous):

actually*

OpenStudy (ybarrap):

then \[(9/8)^{1/2}\]

OpenStudy (anonymous):

thank you everybody